A coin is tossed and a si-sided die numbered 1 through 6 is rolled. Find the probability of tossing a head and then rolling a number greater than 3 The probability of tossing a head and then rolling a number greater than 3 is (Round to three decimal places as needed.)
Find the indicated area under the standard normal curve. To the left of z =-0.91 Click here to view page 1 of the standard normal table, Click here to view page 2 of the standard normal table. The area to the left of z=-0.91 under the standard normal curve is (Round to four decimal places as needed)
Let p be the population proportion for the following condition. Find the point estimates for p and q Of 923 children surveyed, 113 plan to join the armed forces in the future. The point estimate for p, p, is Round to three decimal places as needed.) The point estimate for q. q, is Round to three decimal places as needed.)
Determine whether the following statement is true or false. If it is false, rewrite it as a true statement If two events are independent, P(AB)= P(B). Choose the correct answer below O A. False, if events A and B are independent, then P(BIA) = P(A) O B. False, if events A and B are independent, then P(A and B)-P(A)-P(B). O C. True 0 D. False: if events A and B are independent, then P(A and B) = 0
If two events are mutually exclusive, why is P(A and B)-02 Choose the correct answer below. O A. P(A and B)-0 because A and B are independent. O B. P(A and B)-0 because A and B cannot occur at the same time. O C. P(A and B) 0 because A and B each have the same probability O D. P(A and B) 0 because A and B are complements of each other.
Assume the random variable x is normally distributed with mean μ-60 and standard deviation σ-7 Find the indicated probability P(x> 42) P(x > 42)-□ (Round to four decimal places as needed.)
In a survey of 2255 adults, 731 say they believe in UFOs. Construct a 90% confidence interval for the population proportion of adults who believe in UFOs A 90% confidence interval for the population proportion is Round to three decimal places as needed.) Interpret your results. Choose the correct answer below. 0 A With 90% probability, the population proportion of adults who do not believe in UFOs is between the endpoints of the given confidence inte vaL B C. O D. With 90% confidence it can be said that the population proportion of adults who believe in UFOs is between the endpoints of the given confidence interval With 90% confidence can be said that the sample p portion of adults who believe in UFOs is between the endpoints of the given confidence interval The endpoints of the given confidence interval shows that 90% of adults believe in UFOs.
You roll a six-sided die. Find the probability of each of the following scenarios. (a) Rolling a 6 or a number greater than 3 (b) Rolling a number less than 4 or an even number (c) Rolling a 4 or an odd number (a) P(6 or number > 3)-(Round to three decimal places as needed.) (b) P/1 or 2 or 3 o r 4 or 6) (Round to three decimal places as needed.) (c) P/4 or 1 or 3 or 5) Round to three decimal places as needed.)
able to find the 2-score that corresponds to the cumulative area 0 1190 I the area is not in the table,use the entry closest to the area If the area is hallway between two entries, use the z-score haltway between the cornresponding z-scores z-L」(Type an integer or decimal rounded to three decimal places as needed)
media%2F01a%2F01ac5534-5cfa-46c7-b814-b2
Answer #1

:

Problem 1

Sample space = ( H1,H2,H3,H4,H5,H6,T1,T2,T3,T4,T5,T6)

Possible options : 12

Number of head and number greater than 3 : 3

The required P= 3/12 = 0.250

Problem 2

P( z < -0.91)= 0.1814

Problem 3

p=113/923 =0.122

q=0.878

problem 4

B False. If events are independent, P( A and B) = P(A)*P(B)

Problem 5

D. P( A and )B)=0 because A and B are complements of each other.

Problem 6

Z value for 42, z=(42-50)/7 = -1.14

P( x >42) =P( z > -1.14)

=0.8729

Problem 8

Confidence Interval Estimate for the Proportion

Data

Sample Size

2255

Number of Successes

731

Confidence Level

90%

Intermediate Calculations

Sample Proportion

0.324168514

Z Value

-1.6449

Standard Error of the Proportion

0.0099

Interval Half Width

0.0162

Confidence Interval

Interval Lower Limit

0.3080

Interval Upper Limit

0.3404

90% CI = (0.308, 0.340)

Option B.